# A Smart Mnemonic Sounds like “Glue Tonic”: Mixing LLMs with Student Feedback to Make Mnemonic Learning Stick

### v1

@@ 3.2.2 Observed Preferences
As keyword mnemonics can improve short-term recall Wang et al. (1992), we base observed preferences on the mean turns $t_{X}$ users studying with $m_{X}$ need to correctly answer its card $f_{v}$, a proxy for short-term learning. For each pair $(m_{A},m_{B})$, if $t_{A}<t_{B}$, we call this preference label $y_{learn}$ and set $y_{learn}=\texttt{A}$, as $m_{A}$ helps users learn the definition of $v$ quicker than $m_{B}$ (same for B and tie). $y_{learn}$ is collected automatically as users study.
Figure 4: Screenshot of UI for pairwise comparisons.
@@ 3.3 User Study Details
We deploy mnemonics from $\mathcal{D}_{pref}$ and have 47 students from exam preparation forums, Google ads, and university courses study in our app. We add random quality checks in pairwise comparisons, where one mnemonic is clearly low-quality. All users, including the two users that failed the quality checks, are given \$50. In three months, 45 students gave 2684 preferences for 752 mnemonic pairs and 472 unique pairs (Table 5, details in Appendix A.2). We omit pairs with two or fewer labels so the mean labels per pair is 3.57, following the method of using three or more preferences to curb noise Bai et al. (2022); Ji et al. (2024). Users are referred to by ID.
@@ 4 Preference Analysis
We study the relation of our three preferences (section 3.2) and uncover that students cannot fully predict what helps them learn (section 4.1, section 4.2). This leads us to intuit that our preferences capture complementary facets of mnemonic effectiveness, inspiring the design of our final model that combines all preferences (section 5).
4.1 Are Preference Types Equivalent?
To see if our preference labels capture equivalent information, we compute the agreement of preferences (e.g. $y_{pair}$ vs $y_{rate}$) for the same mnemonics. We exclude labels denoting a tie, focusing instead on labels that show a clear preference towards one mnemonic. Table 1 shows that the expressed preferences $y_{pair}$ and $y_{rate}$ have moderate agreement ($0.675$), aligning with Bansal et al. (2024), who also uncover disagreement in pairwise and rating preferences. But notably, the agreement between expressed and observed preferences is much lower ($0.507$ and $0.407$), meaning that asking students which mnemonics they think are more helpful cannot capture what is truly more helpful for learning.
@@ 4.1 Are Preference Types Equivalent?
Table 1: Raw agreement of preference label types.
Figure 5: Correlation between a user mnemonic rating and the iterations needed for the same user to recall the term while studying with said mnemonic (jittered).
@@ 4.2 Can Users Predict their Learning?
To study preference type disagreement at the student level, we see if a user’s rating of a mnemonic predicts the total turns needed for the user to learn the vocab term linked to that mnemonic. If users can predict this well, we would see a strong negative correlation, with lower ratings indicating more turns needed, but we find little ($r=-0.06$) correlation (Figure 5). Prior work shows that students struggle to identify study strategies that best aid learning McCabe (2011); Yan et al. (2016), and we uncover that students struggle to do the same for study items, further showcasing that expressed preferences cannot fully capture learning outcomes.
@@ 5 Training a Smart-er Mnemonic Model
Our goal of collecting student preference labels for mnemonic pairs is to identify “better” or more effective mnemonics, using this signal to guide Smart’s outputs. But as preference types disagree (section 4), how should we identify more effective mnemonics?
@@ 5.1 Learning Mnemonic Effectiveness
For mnemonic pairs $(m_{A},m_{B})\in\mathcal{D}_{pref}$, we seek to find the more effective mnemonic. We intuit that mnemonic effectiveness is a latent value that can be modeled via feedback from all three of our preferences. Bayesian models are well-suited for this task as they capture annotator noise in feedback more effectively than aggregation Wang et al. (2023a).
Thus, we design a Hierarchical Bayesian model Hoyle et al. (2019) to estimate mnemonic effectiveness. We aim to learn $\text{I\kern-1.49994ptP}(\theta_{A}>\theta_{B})$, the probability mnemonic $m_{A}$ is more effective than mnemonic $m_{B}$. To do so, we individually model mnemonics $m_{A,i}$ and $m_{B,i}$ by latent effectiveness parameters $\theta_{A,i}$ and $\theta_{B,i}$, which are assigned uniform priors:
$\theta_{A,i},\theta_{B,i}$ $\sim\mathrm{Beta}(1,1)$ (4)
@@
For pairwise preferences, let $\mathcal{C}_{i}=\{c_{1},...,c_{n}\}$ be the pairwise ratings of $(m_{A,i},m_{B,i})$, where $c_{i}\in\{\texttt{A},\texttt{B},\texttt{tie}\}$. If $\theta_{A,i}>\theta_{B,i}$, we assume that $\mathcal{C}_{i}$ has more A ratings. To model $\mathcal{C}_{i}$ from $\theta_{A,i}$ and $\theta_{B,i}$, we first compute the sigmoid ($\sigma$) of a linear transform of $\theta_{A,i}$ as the probability $p^{pair}_{A,i}=\text{I\kern-1.49994ptP}(\texttt{A}\in\mathcal{C}_{i})$:
@@
For rating preferences, let $\mathcal{R}_{A,i}=\{r_{1},...,r_{5}\}$ be cumulative counts of Likert ratings for $m_{A,i}$, where $m_{A,i}$ has $r_{j}$ votes less than or equal to rating $j$. We assume an $m_{A,i}$ with higher effectiveness $\theta_{A,i}$ has higher ratings. We model $\mathcal{R}_{A,i}$ as a multinomial distribution, parameterized by a linear transformation of $\theta_{A,i}$ to a 5-length probability distribution:
@@
and same for $\mathcal{R}_{B,i}$. For learning preferences, let $\mathcal{T}_{A,i}=\{t_{1},...,t_{m}\}$ be a distribution of turns users need to recall the term with $m_{A,i}$, where $t_{j}\in\mathbb{Z}^{+}$. We assume an $m_{A,i}$ with higher effectiveness $\theta_{A,i}$ yields less $t_{j}$ needed. $t_{j}$ is the tries until a success, so we model each $t_{j}$ as a Geometric distribution parameterized by a linear transformation of $\theta_{A,i}$:
@@
and same for $\mathcal{T}_{B,i}$. We learn parameters via NUTS Hoffman et al. (2014) for 1000 epochs. Parameters converge across five chains (Appendix B.4), meaning our model consistently estimates effectiveness.
@@ 5.2 Aligning Smart with Student Preferences
We now use the learned effectiveness of mnemonics $(m_{A},m_{B})$ in the preference dataset $\mathcal{D}_{pref}$ to align Smart. Among many methods, we adopt Direct Preference Optimization (Rafailov et al., 2024, DPO), which optimizes on preferences without reward modeling or reinforcement learning. Alternatives like Proximal Policy Optimization Schulman et al. (2017) need extensive parameter tuning and are thus harder to reproduce Huang et al. (2022b).
DPO requires dataset entries with a prompt $x$ and winning/losing outputs $y_{w}$/$y_{l}$, where $y_{w}$/$y_{l}$ are “good”/“bad” outputs for $x$. We set $x$ to the term $v$ in $\mathcal{D}_{pref}$ with its mnemonics $(m_{A},m_{B})$ as outputs. $y_{w}$ is the mnemonic with higher effectiveness ($\theta_{A}$ vs $\theta_{B}$ from section 5.1), and the other is $y_{l}$. With this data, we update our initial model $p_{0}(m|v)$ ($\pi_{0}$ below) to align a better Smart model $p_{dpo}(m|v)$ ($\pi$ below) with DPO, which minimizes the loss $\mathcal{L}_{dpo}$:
$\mathcal{L}_{dpo}=\mathop{-\mathbb{E}}_{\begin{subarray}{c}x,y_{w},y_{l}\\ \sim\mathcal{D}_{pref}\end{subarray}}\Big{[}\ln\sigma\Big{(}\beta\ln\frac{\pi(y_{w}|x)}{\pi_{0}(y_{w}|x)}-\beta\ln\frac{\pi(y_{l}|x)}{\pi_{0}(y_{l}|x)}\Big{)}\Big{]}.$ (16)
@@ 6 How Smart Are Smart’s Mnemonics?
Given the costs of human pairwise evaluations, we adopt a standard practice of having GPT-4 judge which of two model-created mnemonics is higher quality Chiang et al. (2023); Liu et al. (2023). GPT-4 shows 80% agreement with users on 200 held-out mnemonic pairs (Appendix B.3), near the 81% human agreement in MT-bench Zheng et al. (2024), so GPT-4 agrees with user pairwise mnemonic ratings. To curb position bias Wang et al. (2023b), we compare mnemonics in both orders, only marking that one model wins if GPT-4 picks the model’s mnemonic in both orders, otherwise marking a tie.
We first use GPT-4 to compare mnemonic quality of Smart versions (section 6.1, section 6.2). We then have mnemonic experts evaluate Smart’s mnemonics to inform future work (section 6.3). We also present examples of Smart’s mnemonics in Appendix D.4.
@@ 6.1 Ablation Study
We ablate Smart (Figure 1) to verify each step improves mnemonic quality. Our fine-tuned model $p_{0}(m|v)$ creates higher-quality mnemonics versus few-shot LLaMA $p_{fs}(m|v)$ prompted using the 10 $\mathcal{D}_{ft}$ examples with the highest latent quality (section 2.2), and same for $p_{dpo}(m|v)$ versus $p_{0}(m|v)$ (Table 2). Both of the steps improve pairwise mnemonic quality, confirming DPO can align LLMs with student preferences to enhance LLM outputs in education.
@@
Table 2: GPT judgement of our ablations. Significantly better models (Binomial, $p<0.005$) are bold with *.
@@
Table 3: GPT judgement of DPO models. Significantly better models (Binomial, $p<0.005$) are bold with *.
6.2 DPO with Multiple Preference Labels
We study the effects of using DPO with MPLs for pairwise mnemonic quality through two questions:
Q1—Do MPLs harm pairwise metrics? One concern of using MPLs for DPO is that the model will produce lower-quality mnemonics compared to a model using pairwise labels, as the latter optimizes just on the evaluation metric. To test this, we take a subset of preference data $\mathcal{D}_{pair}\subset\mathcal{D}_{pref}$ [^2] To have enough label disagreement, $\mathcal{D}_{pref}$ in this analysis also adds mnemonics with two labels (see Appendix A.3). with the pairwise preference $y_{pair}$. We train two DPO models on $\mathcal{D}_{pair}$ when $y_{pair}\neq\texttt{tie}$: $p_{bayes}(m|v)$, training on the Bayesian label $y_{bayes}$ from section 5.1, and $p_{pair}(m|v)$, training on $y_{pair}$. Despite 20% disagreement in $y_{pair}$ and $y_{bayes}$ on the winning mnemonic, the models are judged to generate equal-quality mnemonics on $\mathcal{V}_{test}$ (Table 3, top). Thus, DPO training with MPLs does not always degrade LLM output quality on singular preference metrics.
Q2—Can MPLs augment data? If we train DPO with just $y_{pair}$, we must omit data where $y_{pair}=\texttt{tie}$ or no $y_{pair}$ exists. While we could gain more $y_{pair}$ labels with another user study, we see if $y_{bayes}$ can resolve missing or tied $y_{pair}$ labels using other preferences to elect winning mnemonics, augmenting data without more pairwise labels. We compare $p_{pair}(m|v)$, which trains on 348 $y_{pair}$ labels, to our full model $p_{dpo}(m|v)$, which trains on 117 extra pairs when $y_{bayes}$ breaks a tie in $y_{pair}$ and 12 extra pairs with no $y_{pair}$ label. $p_{dpo}(m|v)$ generates significantly better ($p<0.005$) mnemonics than $p_{pair}(m|v)$ (Table 3, bottom), meaning that MPLs can effectively augment DPO training data over pairwise preferences for LLM output quality gains.
@@ 6.3 Qualitative Evaluation
For a detailed evaluation, we have two mnemonic researchers assess our mnemonics, split into keyword and explanation quality. For keyword quality, we ask two yes/no questions: 1) Does the keyword sound like the term? (Phonetic Similarity); 2) Is the keyword simpler [^3] A keyword is simpler if a user not knowing the term could likely know the keyword (torpor is not simpler than torpid). than the term? (Simplicity). For explanation quality, we rate mnemonic explanations out of five on: 1) Clarity: Ease of understanding; 2) Strength: The obviousness of the explanation’s association of the keyword and the term Hall et al. (1981); 3) Imageability: The ability to evoke mental imagery Campos et al. (2011). We argue clarity and strength are most important for explanation quality, ensuring students can understand and create strong memory links from terms to keywords. We use imageability as it can affect memory Groninger (1971), but an imageable explanation is still unmemorable if unclear or low strength. [^4] For example, Ben Franklin for benevolent in Figure 1.
Figure 6: Expert qualitative evaluation of mnemonics.
@@
We assess mnemonic keyword and explanation quality for 50 $\mathcal{V}_{test}$ terms created by a professional writer from Upwork, 10-shot GPT-4, and Smart ($p_{dpo}(m|v)$). We also compare Smart’s keyword quality to Transphoner Savva et al. (2014), a SOTA mnemonic keyword extractor. Transphoner is hard to reproduce, so we use 50 terms and Transphoner outputs released by the authors for this comparison.
@@
Keyword Quality: Smart has slightly better keywords than Transphoner (higher simplicity, equal PS), meaning LLMs are strong alternatives for keyword extractors (Figure 6, left). Smart also has much simpler keywords than GPT-4, but with lower PS (Figure 6, middle). Suvarna et al. (2024) finds a large gap (37% accuracy) when prompting LLaMA-2 and GPT-4 for rhyme generation which also tests PS; our gap is just 8%, so fine-tuning and user feedback help LLMs address phonetic weaknesses. Our writer bests both models in keyword quality, with large simplicity gaps. Thus, models that simplify text with LLMs (summarizers, topic models) may benefit from explicit feedback on word simplicity.
@@
Explanation Quality: Smart matches the explanations of GPT-4 (lower clarity, higher strength), meaning student feedback enables smaller LLMs like Smart to compete with SOTA LLMs (Figure 6, right). Our writer again has the best explanations, especially in imageability over Smart, as Smart is not guided toward imageable mnemonics. Imageability does not entail memorability, but it may be useful for tasks in memorability (story-telling, advertising) to use imageability feedback, as even GPT-4 shows a weakness in imageability.
@@
Takeaway: Training LLMs with student feedback results in mnemonics with keyword quality matching SOTA extractive methods while also generating explanations. It also allows Smart, a smaller LLM, to compete with GPT-4, so using such feedback with GPT-4 could yield even better mnemonics. Lastly, our expert writes better mnemonics than LLMs, specifically in keyword simplicity and imageability. This motivates mnemonic generation as a challenging task and gives insights into feedback that could be used to guide LLMs in similar tasks.

### v2

@@ 3.2.2 Observed Preferences
As keyword mnemonics can improve short-term recall Wang et al. (1992), we base observed preferences on the mean turns $t_{X}$ users studying with $m_{X}$ need to correctly answer its flashcard $f_{v}$, as calculated in section 3.2. This measure forms a proxy for short-term learning. For a given pair $(m_{A},m_{B})$, if $t_{A}<t_{B}$, we call this preference label $y_{learn}$ and set $y_{learn}=\texttt{A}$, as $m_{A}$ helps users learn the definition of $v$ quicker than $m_{B}$ (same for B and tie). $y_{learn}$ is collected automatically as users study.
@@ 3.3 User Study Details
We deploy all mnemonics from $\mathcal{D}_{pref}$ and have 47 English-speaking students from exam preparation forums, Google ads, and university courses study in our app. To filter noisy annotators, we add random quality checks in pairwise comparisons, where one mnemonic is clearly low-quality. If any user picks the low-quality mnemonic in the pairwise comparison, their preferences are omitted for analysis and model training. All users, including the two users who failed our quality checks, are awarded \$50. In three months, 45 students gave us 2684 preferences for 752 mnemonic pairs and 472 unique pairs (Table 5, details in Appendix A.2). We omit pairs with two or fewer labels so the mean labels per pair is 3.57, following the method of using three or more preferences to curb noise Bai et al. (2022); Ji et al. (2024). Users are referred to just by numerical ID.
@@ 4 Preference Analysis
We study the relation of our preference types (section 3.2) and uncover that students cannot fully predict what aids learning (section 4.1, section 4.2). Thus, we conclude that a mnemonic’s overall helpfulness cannot be captured by just one preference type, inspiring the design of our final model that combines all preferences (section 5).
@@
Table 1: Raw agreement of preference types. Expressed preferences $y_{pair}$ and $y_{rate}$ have some disagreement, but agreement between expressed and observed preferences ($y_{rate}$ vs $y_{learn}$ and $y_{pair}$ vs $y_{learn}$) is even lower.
4.1 Are Preference Types Equivalent?
To see if our preference labels capture equivalent information, we compute the agreement of preferences (e.g. $y_{pair}$ vs $y_{rate}$) for the same mnemonics. We exclude labels denoting a tie, focusing instead on labels that show a clear preference towards one mnemonic. Table 1 shows that the expressed preferences $y_{pair}$ and $y_{rate}$ have moderate agreement ($0.675$), aligning with Bansal et al. (2024), who also uncover disagreement in pairwise and rating preferences. But notably, the agreement between expressed and observed preferences is much lower ($0.507$ and $0.407$), so asking students which outputs they think are more helpful does not always capture what is truly more helpful for learning.
Figure 5: Correlation between user mnemonic ratings and turns needed for the same user to recall the term when studying with said mnemonic (jittered). Users cannot predict which mnemonics will best help them learn.
@@ 4.2 Can Users Predict their Learning?
To study preference type disagreement at the student level, we see if a user’s rating of a mnemonic predicts the total turns needed for the user to learn the vocab term linked to that mnemonic. If users can predict this well, we would see a strong negative correlation, with lower ratings indicating more turns needed, but we find little ($r=-0.06$) correlation (Figure 5). Prior work shows that students struggle to identify the study strategies that best aid learning McCabe (2011); Yan et al. (2016), and we uncover that students struggle to do the same for study items, further showcasing that expressed preferences cannot fully capture learning outcomes.
@@ 5 Training a Smart-er Mnemonic Model
Our goal of collecting student preference labels for mnemonic pairs is to identify more helpful or effective mnemonics, using this signal to guide Smart’s outputs. But as preference types disagree (section 4), how should we identify more effective mnemonics?
@@ 5.1 Learning Mnemonic Effectiveness
For mnemonic pairs $(m_{A},m_{B})\in\mathcal{D}_{pref}$, we seek to find the more effective mnemonic. We intuit that mnemonic effectiveness is a latent value that can be modeled via feedback from all three of our preferences. Bayesian models are well-suited for this task as they capture annotator noise in feedback more effectively than aggregation Wang et al. (2023).
Thus, we design a Hierarchical Bayesian model Gelman and Hill (2006) to estimate mnemonic effectiveness. We seek to learn $\text{I\kern-1.49994ptP}(\theta_{A}>\theta_{B})$, the probability mnemonic $m_{A}$ is more effective than mnemonic $m_{B}$. To do so, we model mnemonics $m_{A,i}$ and $m_{B,i}$ by latent effectiveness parameters $\theta_{A,i}$ and $\theta_{B,i}$, which are assigned uniform priors:
$\theta_{A,i},\theta_{B,i}$ $\sim\mathrm{Beta}(1,1).$ (4)
@@
For pairwise preferences, let $\mathcal{C}_{i}=\{c_{1},...,c_{n}\}$ be the pairwise ratings of $(m_{A,i},m_{B,i})$, where $c_{i}\in\{\texttt{A},\texttt{B},\texttt{tie}\}$. If $\theta_{A,i}>\theta_{B,i}$, we assume $\mathcal{C}_{i}$ has more A preferences. To model $\mathcal{C}_{i}$ from $\theta_{A,i}$ and $\theta_{B,i}$, we compute the sigmoid ($\sigma$) of a linear transform of $\theta_{A,i}$ as the probability $p^{pair}_{A,i}=\text{I\kern-1.49994ptP}(\texttt{A}\in\mathcal{C}_{i})$:
@@
For rating preferences, let $\mathcal{R}_{A,i}=\{r_{1},...,r_{5}\}$ be cumulative counts of Likert ratings for $m_{A,i}$, where $m_{A,i}$ has $r_{j}$ votes less than or equal to rating $j$. We assume $m_{A,i}$ with higher effectiveness $\theta_{A,i}$ has higher ratings. We model $\mathcal{R}_{A,i}$ as a multinomial distribution, parameterized by a linear transformation of $\theta_{A,i}$ to a 5-length probability distribution:
@@
and do the same for $\mathcal{R}_{B,i}$.
For learning preferences, $\mathcal{T}_{A,i}=\{t_{1},...,t_{m}\}$ is the distribution of turns users need to recall the term with $m_{A,i}$, where $t_{j}\in\mathbb{Z}^{+}$. We assume $m_{A,i}$ with higher effectiveness $\theta_{A,i}$ yields fewer $t_{j}$ needed. Every turn count $t_{j}$ is the tries until a success, so we model each $t_{j}$ as a Geometric distribution parameterized by a linear transformation of $\theta_{A,i}$:
@@
and same for $\mathcal{T}_{B,i}$. We learn all variables via NUTS Hoffman et al. (2014) for 1000 epochs. Parameters converge across five chains (Appendix B.4), meaning our model consistently estimates effectiveness.
@@ 5.2 Aligning Smart with Student Preferences
We now use the learned effectiveness of mnemonics $(m_{A},m_{B})$ in the preference dataset $\mathcal{D}_{pref}$ to align Smart. Among many alignment methods, we adopt Direct Preference Optimization (Rafailov et al., 2024, DPO), which tunes LLMs to preferences without requiring explicit reward modeling or reinforcement learning steps. Alternatives like Proximal Policy Optimization Schulman et al. (2017) need extensive parameter tuning and are thus harder to reproduce Huang et al. (2022b).
DPO requires dataset entries with a prompt $x$ and winning/losing outputs $y_{w}$/$y_{l}$, where $y_{w}$/$y_{l}$ are “good”/“bad” outputs for $x$. We set $x$ to the term $v$ in $\mathcal{D}_{pref}$ with its mnemonics $(m_{A},m_{B})$ as outputs. The mnemonic with higher effectiveness ($\theta_{A}$ vs $\theta_{B}$ from section 5.1) is $y_{w}$, and the other is $y_{l}$. With this data, we update our initial model $p_{0}(m\,|\,v)$ ($\pi_{0}$ below) to align a better Smart model $p_{dpo}(m\,|\,v)$ ($\pi$ below) with DPO, which minimizes the loss $\mathcal{L}_{dpo}$:
$\mathcal{L}_{dpo}=\mathop{-\mathbb{E}}_{\begin{subarray}{c}x,y_{w},y_{l}\\ \sim\mathcal{D}_{pref}\end{subarray}}\left[\ln\sigma\left(\beta\ln\frac{\pi(y_{w}|x)}{\pi_{0}(y_{w}|x)}-\beta\ln\frac{\pi(y_{l}|x)}{\pi_{0}(y_{l}|x)}\right)\right].$ (16)
@@ 6 How Smart Are Smart’s Mnemonics?
Given the costs of human pairwise evaluations, we adopt a common practice having GPT-4 judge which of two model-created mnemonics is higher quality Chiang et al. (2023); Liu et al. (2023). GPT-4 has 80% agreement with users (Appendix B.3) on 200 held-out mnemonic pairs, near the 81% human agreement in MT-bench Zheng et al. (2024), so GPT-4 agrees with user pairwise mnemonic ratings. To curb position bias Wang et al. (2024), we compare mnemonics in both orders, only marking that one model wins if GPT-4 picks the model’s mnemonic in both orders, otherwise marking a tie.
We first use GPT-4 to compare mnemonic quality of Smart ablations (section 6.1, section 6.2). We then have mnemonic experts evaluate Smart’s mnemonics to inform future work (section 6.3). We also present examples of Smart’s mnemonics in Appendix D.4.
@@ 6.1 Ablation Study
We ablate Smart (Figure 1) to verify that both fine-tuning and DPO improve mnemonic quality. Our fine-tuned model $p_{0}(m\,|\,v)$ generates higher-quality mnemonics versus the few-shot LLaMA model $p_{fs}(m\,|\,v)$ prompted using the ten $\mathcal{D}_{ft}$ examples with the highest latent quality (section 2.2), and same for $p_{dpo}(m\,|\,v)$ versus $p_{0}(m\,|\,v)$ (Table 2). Both of the steps improve pairwise mnemonic quality, confirming DPO can align LLMs with student preferences to enhance LLM outputs in education.
6.2 DPO with Multiple Preference Labels
We investigate the effectiveness of training with DPO using MPLs for pairwise mnemonic quality through two research questions, outlined below:
Q1—Do MPLs harm pairwise metrics? One concern of optimizing on MPLs with DPO is that the model will produce lower-quality mnemonics compared to a model using pairwise labels, as the latter optimizes just on the evaluation metric. To test this, we first select a subset of preference data $\mathcal{D}_{pair}\subset\mathcal{D}_{pref}$ [^2] To have enough label disagreement, $\mathcal{D}_{pref}$ in this analysis also adds mnemonics with two labels (see Appendix A.3). with the pairwise preference $y_{pair}$. We then train two DPO models on $\mathcal{D}_{pair}$ when $y_{pair}\neq\texttt{tie}$: $p_{bayes}(m\,|\,v)$, training on the Bayesian label $y_{bayes}$ from section 5.1, and $p_{pair}(m\,|\,v)$, training on $y_{pair}$. Despite having 20% disagreement in $y_{pair}$ and $y_{bayes}$ on the winning mnemonic, the two models are judged to generate mnemonics with equal quality on $\mathcal{V}_{test}$ (Table 3, top). Thus, DPO training with MPLs does not always degrade LLM output quality on singular preference metrics.
@@ 6.2 DPO with Multiple Preference Labels
Table 2: GPT judgement of our ablations. Significantly better models (Binomial, $p<0.005$) are bold with *. Our fine-tuning and DPO steps both improve Smart.
@@
Table 3: GPT judgement of DPO models. Significantly better models (Binomial, $p<0.005$) are bold with *. Multiple preferences can break ties in singular preferences for mnemonic quality gains (bottom row).
Q2—Can MPLs augment data? If we train DPO with just $y_{pair}$, we must discard preference data when $y_{pair}=\texttt{tie}$ or no $y_{pair}$ exists. While we could collect more $y_{pair}$ labels with another user study, we investigate whether $y_{bayes}$ can directly resolve the missing or tied $y_{pair}$ labels using the other preferences to elect winning mnemonics, augmenting our training data without collecting any more pairwise comparison data. We compare $p_{pair}(m\,|\,v)$, which trains on 348 $y_{pair}$ labels, to our full model $p_{dpo}(m\,|\,v)$, which trains on 117 extra pairs when $y_{bayes}$ breaks a tie in $y_{pair}$ and twelve extra pairs without $y_{pair}$ labels. $p_{dpo}(m\,|\,v)$ has significantly better ($p<0.005$) mnemonics than $p_{pair}(m\,|\,v)$ (Table 3, bottom), meaning that MPLs can effectively augment DPO training data over pairwise preferences for output quality gains.
@@ 6.3 Qualitative Evaluation
For a detailed evaluation, we have two mnemonic researchers assess our mnemonics, split into keyword and explanation quality. For keyword quality, we ask two yes/no questions: 1) Does the keyword sound like the term? (Phonetic Similarity); 2) Is the keyword simpler [^3] A keyword is simpler if a user not knowing the term could likely know the keyword (torpor is not simpler than torpid). than the term? (Simplicity). For explanation quality, we rate mnemonic explanations out of five on: 1) Clarity: Ease of understanding; 2) Strength: The obviousness of the explanation’s association of the keyword and the term Hall et al. (1981); 3) Imageability: The ability to evoke mental imagery Campos et al. (2011). We argue clarity and strength are most important for explanation quality, ensuring students can understand and create strong memory links from terms to keywords. We use imageability as it can affect memory Groninger (1971), but an imageable explanation is still unmemorable if it is unclear or low strength. [^4] For example, Ben Franklin for benevolent in Figure 1.
@@
We assess mnemonic keyword and explanation quality for 50 $\mathcal{V}_{test}$ terms created by a professional writer from Upwork, 10-shot GPT-4, and Smart, $p_{dpo}(m\,|\,v)$. We also compare Smart’s keyword quality to Transphoner Savva et al. (2014), a SOTA mnemonic keyword extractor. Transphoner is hard to reproduce, so we use 50 terms and Transphoner outputs released by the authors for this comparison.
@@
Keyword Quality: Smart has slightly better keywords than Transphoner (higher simplicity, equal PS), meaning LLMs are strong alternatives for keyword extractors (Figure 6, left). Smart also produces much simpler keywords than GPT-4, but with lower PS (Figure 6, middle). Suvarna et al. (2024) reveals a large gap (37% accuracy) when prompting LLaMA-2 13B and GPT-4 for rhyme generation which also assesses PS; our gap is just 8%, so fine-tuning and human feedback can help LLMs address phonetic weaknesses. Lastly, our human writer surpasses both models in keyword quality, with large simplicity gaps. Thus, systems that simplify text with LLMs (summarizers, topic models) may benefit from explicit feedback on word simplicity.
Figure 6: Expert qualitative evaluation of mnemonics. Smart matches GPT-4 at much lower deployment costs, but our human writer largely surpasses all models, especially in keyword simplicity and imageability.
@@
Explanation Quality: Smart matches GPT-4’s explanations (lower clarity, but higher strength), meaning student feedback lets smaller LLMs like Smart compete with large SOTA LLMs (Figure 6, right). Our writer again has the best explanations, especially in imageability over Smart, as Smart is not explicitly guided toward imageable outputs. Imageability does not entail memorability, but it may be useful for memorability tasks (story-telling, advertising) to use imageability feedback, as even GPT-4 shows a clear weakness in imageability.
@@
Takeaway: Training LLMs with student feedback results in mnemonics with keyword quality matching SOTA extractive methods while also generating explanations. It also allows Smart, a smaller LLM, to compete with GPT-4. This is a significant feat, as Smart can provide students with mnemonics as effective as GPT-4 at much lower deployment costs—an attractive feature for educators looking to implement LLM educational tools. Further, since GPT-4 is stronger than LLaMA, using student feedback with GPT-4 could yield even better mnemonics. Lastly, our expert writes better mnemonics than LLMs, specifically in keyword simplicity and imageability. This motivates mnemonic generation as a challenging task and gives insights into feedback that could be used to guide LLMs in similar tasks.